You have 100 marbles, numbered 1-100. If you pick one, what is the probability it will either be odd or greater than 80?
Answer by richard1234(5390)

P(odd or >80) = P(odd) + P(>80) - P(odd U >80) = 50/100 + 20/100 - 10/100 = 60/100 = 3/5